# SWE-Pruner: Self-Adaptive Context Pruning for Coding Agents

### v1

@@ 1 Introduction
Large Language Models (LLMs) have rapidly progressed in software engineering tasks, from simple code understanding [chen2021evaluating, li2023starcoder, yang2025elaboration, hu2025flowmaltrans] to interactive agents capable of navigating repositories, running tests, and submitting patches end-to-end [chen2025swe, li2025swe, xiao2025improving]. Recent coding agents like Claude Code [claude_code] and Gemini CLI [gemini_cli] have augmented LLMs with sophisticated toolchains including terminals, editors, and file search, enabling multi-step reasoning over complex codebases.
Despite this progress, a critical bottleneck remains: the accumulation of context length within the constrained window of LLMs. Navigating real-world software repositories confronts agents with a massive “Context Wall.” Although long-context models have emerged [achiam2023gpt, team2024gemini], blindly ingesting large volumes of code incurs prohibitive inference costs [jiang2023llmlingua] and introduces severe noise, leading to attention dilution and hallucinations [liu2023lost, li2023loogle, wang2025position].
@@
We evaluate SWE-Pruner across multi-turn agent tasks (SWE-Bench Verified [jimenez2024swe] and SWE-QA [peng2025swe]) and single-turn long-context understanding tasks (Long Code Completion [guo2023longcoder] and Long Code QA [rando2025longcodebench]). Across models and benchmarks, SWE-Pruner consistently delivers substantial efficiency gains with minimal performance degradation. Beyond token savings (e.g., 39% reduction on SWE-Bench Verified with Claude Sonnet 4.5, as illustrated in Figure˜1), the focused context also improves agent decision quality, reducing interaction rounds by up to 26% through more decisive reasoning and fewer redundant exploratory actions.
@@ 5.1 Performance on Multi-Turn Tasks
Table 1: Results on SWE-Bench Verified. SWE-Pruner reduces token consumption by 23–38% while maintaining comparable success rates.
@@
We integrate SWE-Pruner into the Mini SWE Agent [yang2024sweagent] framework for SWE-Bench Verified [jimenez2024swe] and the OpenHands [openhands] framework for SWE-QA [peng2025swe], evaluating with different backbone models. On SWE-Bench Verified, SWE-Pruner achieves substantial token reductions of 23–38% across models while maintaining nearly identical success rates (less than 1% degradation). Notably, interaction rounds decrease by 18–26%. By filtering redundant information while preserving task-relevant code, SWE-Pruner enables agents to locate issues more precisely and make more decisive decisions, thereby reducing repeated exploratory file reads and completing tasks earlier. We conduct in-depth case studies in Appendix˜I to illustrate these behavioral differences, showing how context pruning transforms both failure-to-success scenarios (with 83.3% token reduction) and successful trajectories (with 30.2% reduction in peak prompt length). On SWE-QA, similar efficiency gains emerge with 29–54% token reduction across repositories. Interestingly, we observe that GLM-4.6 exhibits increased rounds (29–41% more) while Claude Sonnet 4.5 maintains similar round counts with minor variations (within 3%). Through trajectory analysis, we find that after pruning, GLM tends to explore more files before formulating answers, suggesting a more conservative reasoning strategy when presented with focused context. Nevertheless, the overall token consumption remains substantially lower, demonstrating that SWE-Pruner maintains efficiency advantages regardless of the agent’s exploration patterns. These model-agnostic efficiency gains directly translate to proportional cost savings. Detailed analysis is provided in Appendix˜E.
@@ 6 Related Works
Prompt compression has been extensively studied for both natural language and code. Token-level pruning methods such as LLMLingua [jiang2023llmlingua, jiang2024longllmlingua], Selective Context [li2023compressing], and AttentionRAG [fang2025attentionrag], along with embedding/retrieval-based approaches like RAG [lewis2020retrieval], XRAG [cheng2024xrag], and repo-level retrieval for code completion [zhang2023repocoder], can reduce prompt length on single-round tasks but often fail to preserve syntactic structures essential for code correctness [shi2025longcodezip]. Code-specific methods including DietCode [zhang2022diet], SlimCode [wang2024natural], and LongCodeZip [shi2025longcodezip] address structural concerns but are primarily evaluated on single-round proxy tasks (e.g., code search, completion) rather than multi-round agentic workflows. Moreover, these approaches typically apply static, content-only compression that lacks contextual awareness, leading to fixed behaviors such as indiscriminately removing comments [yang2024less] regardless of task requirements. In contrast, SWE-Pruner performs line-level pruning with dynamic, query-conditioned thresholding that adapts to the agent’s current task stage, enabling context-aware compression validated on end-to-end benchmarks like SWE-Bench [jimenez2024swe] without repository-specific retraining.
@@
Modern coding LLMs support context windows of 128k tokens or more [achiam2023gpt, team2024gemini], yet they remain insufficient for large-scale codebases and suffer from documented performance degradation on long contexts [liu2023lost, laban2025llms, li2023loogle]. Agentic systems [yao2023react, wei2022chain] for software engineering [yang2024sweagent, wang2024openhands, xia2024agentless, bouzenia2024repairagent, qin2024agentfl, liu2024large] face critical challenges in managing multi-turn interaction contexts. In NLP domains, trajectory management approaches employ LLM-based summarization when contexts overflow [cursor, claude_code], fixed-size truncation [gao2025trae], or simple observation masking [lindenbauer2025complexity]. For long-horizon scenarios, recent frameworks such as SUPO [lu2025supo] and FoldGRPO [sun2025contextfolding] learn to manage history through reinforcement learning, while others like COMPASS [wan2025compass], ACON [kang2025acon], AgentDiet [xiao2025improving], and AgentFold [ye2025agentfold] introduce hierarchical oversight or proactive folding policies. These methods primarily optimize prior interaction histories, whereas SWE-Pruner serves as lightweight middleware at the agent–environment boundary to prune the environment’s observation (i.e., file content); it is thus orthogonal to and can be seamlessly combined with such learned history managers.
@@ Appendix C Training Dataset for the Neural Skimmer
We synthesize queries and line-level labels for 200,000 sampled code snippets from 195,370 files across 5,945 repositories. For each code snippet $C$, we employ Qwen3-Coder-30B-A3B-Instruct [bai2023qwen] as the teacher LLM to generate task-oriented queries and corresponding line-level retention masks. Query generation employs a temperature of 0.7 and top-p sampling with $p=0.9$ to balance diversity and coherence. To ensure representativeness, we randomly sample across all nine task types, three snippet length levels (short, medium, long), and three relevance levels (low, medium, high). Following initial generation, we apply a rigorous quality control mechanism using Qwen3-Next-80B-A3B-Thinking as an LLM-as-a-Judge filter [wang2025can, liu2023g, song2024finesure, he2025code]. This judge model evaluates the reasoning quality, annotation consistency, and task alignment of each sample, retaining approximately 1/6 of candidates that meet high-quality standards. The complete generation prompts and judge criteria are detailed in Appendix˜J.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have rapidly progressed in software engineering tasks, from simple code understanding [chen2021evaluating, li2023starcoder, yang2025elaboration, hu2025flowmaltrans, zeng2026codesumm] to interactive agents capable of navigating repositories, running tests, and submitting patches end-to-end [shi2024code, chen2025swe, li2025swe, xiao2025improving, wang2026effiskill]. Recent coding agents like Claude Code [claude_code] and Gemini CLI [gemini_cli] have augmented LLMs with sophisticated toolchains including terminals, editors, and file search, enabling multi-step reasoning over complex codebases.
Despite this progress, a critical bottleneck remains: the accumulation of context length within the constrained window of LLMs. Navigating real-world software repositories confronts agents with a massive “Context Wall.” Although long-context models have emerged [achiam2023gpt, team2024gemini], blindly ingesting large volumes of code incurs prohibitive inference costs [jiang2023llmlingua] and introduces severe noise, leading to attention dilution and hallucinations [liu2023lost, li2023loogle, wang2025position, wang2026fasa].
@@
We evaluate SWE-Pruner across multi-turn agent tasks (SWE-Bench Verified [jimenez2024swe] and SWE-QA [peng2025swe]) and single-turn long-context understanding tasks (Long Code Completion [guo2023longcoder] and Long Code QA [rando2025longcodebench]). Across models and benchmarks, SWE-Pruner consistently delivers substantial efficiency gains while maintaining or even improving task performance. Beyond token savings (e.g., 39% reduction on SWE-Bench Verified with Claude Sonnet 4.5, as illustrated in Figure˜1), the focused context also improves agent decision quality, reducing interaction rounds by up to 26% through more decisive reasoning and fewer redundant exploratory actions.
@@ 5.1 Performance on Multi-Turn Tasks
Table 1: Results on SWE-Bench Verified. SWE-Pruner reduces token consumption by 23–38% while improving success rates by 1.2–1.4 percentage points.
@@
We integrate SWE-Pruner into the Mini SWE Agent [yang2024sweagent] framework for SWE-Bench Verified [jimenez2024swe] and the OpenHands [openhands] framework for SWE-QA [peng2025swe], evaluating with different backbone models. On SWE-Bench Verified, SWE-Pruner achieves substantial token reductions of 23–38% across models while even improving success rates (by 1.2–1.4 percentage points). Notably, interaction rounds decrease by 18–26%. By filtering redundant information while preserving task-relevant code, SWE-Pruner enables agents to locate issues more precisely and make more decisive decisions, thereby reducing repeated reads of exploratory files and completing tasks earlier. We conduct in-depth case studies in Appendix˜I to illustrate these behavioral differences, showing how context pruning transforms both failure-to-success scenarios (with 83.3% token reduction) and successful trajectories (with 30.2% reduction in peak prompt length). On SWE-QA, similar efficiency gains emerge with 29–54% token reduction across repositories. Interestingly, we observe that GLM-4.6 exhibits increased rounds (29–41% more) while Claude Sonnet 4.5 maintains similar round counts with minor variations (within 3%). Through trajectory analysis, we find that after pruning, GLM tends to explore more files before formulating answers, suggesting a more conservative reasoning strategy when presented with focused context. Nevertheless, the overall token consumption remains substantially lower, demonstrating that SWE-Pruner maintains efficiency advantages regardless of the agent’s exploration patterns. These model-agnostic efficiency gains directly translate to proportional cost savings. Detailed analysis is provided in Appendix˜E.
@@ 6 Related Works
Prompt compression has been extensively studied for both natural language and code. Token-level pruning methods such as LLMLingua [jiang2023llmlingua, jiang2024longllmlingua], Selective Context [li2023compressing], and AttentionRAG [fang2025attentionrag], along with embedding/retrieval-based approaches like RAG [lewis2020retrieval, cheng2026resolvingrobustnessprecisiontradeofffinancial, shi2026reasoning], XRAG [cheng2024xrag], and repo-level retrieval for code completion [zhang2023repocoder, lai2026transformers], can reduce prompt length on single-round tasks but often fail to preserve syntactic structures essential for code correctness [shi2025longcodezip]. Code-specific methods including DietCode [zhang2022diet], SlimCode [wang2024natural], LongCodeZip [shi2025longcodezip], CodeOCR [shi2026codeocr], and context inlining [hu2026line, zeng2025pruning] address structural concerns but are primarily evaluated on single-round proxy tasks (e.g., code search, completion) rather than multi-round agentic workflows. Moreover, these approaches typically apply static, content-only compression that lacks contextual awareness, leading to fixed behaviors such as indiscriminately removing comments [yang2024less] regardless of task requirements. In contrast, SWE-Pruner performs line-level pruning with dynamic, query-conditioned thresholding that adapts to the agent’s current task stage, enabling context-aware compression validated on end-to-end benchmarks like SWE-Bench [jimenez2024swe] without repository-specific retraining.
@@
Modern coding LLMs support context windows of 128k tokens or more [achiam2023gpt, team2024gemini], yet they remain insufficient for large-scale codebases and suffer from documented performance degradation on long contexts [liu2023lost, laban2025llms, li2023loogle]. Agentic systems [yao2023react, wei2022chain] for software engineering [yang2024sweagent, wang2024openhands, xia2024agentless, bouzenia2024repairagent, qin2024agentfl, liu2024large, fang2025lastingbench, zhao2026dllm] face critical challenges in managing multi-turn interaction contexts. In NLP domains, trajectory management approaches employ LLM-based summarization when contexts overflow [cursor, claude_code], fixed-size truncation [gao2025trae], or simple observation masking [lindenbauer2025complexity]. For long-horizon scenarios, recent frameworks such as SUPO [lu2025supo] and FoldGRPO [sun2025contextfolding] learn to manage history through reinforcement learning, while others like COMPASS [wan2025compass], ACON [kang2025acon], AgentDiet [xiao2025improving], and AgentFold [ye2025agentfold] introduce hierarchical oversight or proactive folding policies. These methods primarily optimize prior interaction histories, whereas SWE-Pruner serves as lightweight middleware at the agent–environment boundary to prune the environment’s observation (i.e., file content); it is thus orthogonal to and can be seamlessly combined with such learned history managers.
@@ 7 Conclusion
Acknowledgments
This paper is supported by the National Key Research and Development Program of China (Grant No. 2023YFB4503802) and the Natural Science Foundation of Shanghai (Grant No. 25ZR1401175).
@@ Appendix C Training Dataset for the Neural Skimmer
We synthesize queries and line-level labels for 200,000 sampled code snippets from 195,370 files across 5,945 repositories. For each code snippet $C$, we employ Qwen3-Coder-30B-A3B-Instruct [bai2023qwen] as the teacher LLM to generate task-oriented queries and corresponding line-level retention masks. Query generation employs a temperature of 0.7 and top-p sampling with $p=0.9$ to balance diversity and coherence. To ensure representativeness, we randomly sample across all nine task types, three snippet length levels (short, medium, long), and three relevance levels (low, medium, high). Following initial generation, we apply a rigorous quality control mechanism using Qwen3-Next-80B-A3B-Thinking as an LLM-as-a-Judge filter [wang2025can, liu2023g, song2024finesure, he2025code, lan2025f2bench, lan2025mcbe]. This judge model evaluates the reasoning quality, annotation consistency, and task alignment of each sample, retaining approximately 1/6 of candidates that meet high-quality standards. The complete generation prompts and judge criteria are detailed in Appendix˜J.